# EmbodiedOcc++: Boosting Embodied 3D Occupancy Prediction with Plane Regularization and Uncertainty Sampler

### v1

@@ Abstract.
3D occupancy prediction, 3D Gaussian Splatting, Online scene understanding
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
Figure 1. The framework of EmbodiedOcc++. Given monocular RGB inputs, our EmbodiedOcc++ boosts indoor 3D occupancy prediction with plane regularization and uncertainty sampler. We design a Geometry-guided Refinement Module that regularizes Gaussian updates primarily along the tangent plane of surface normal. It determines the adaptive regularization weight using curvature-based and depth-based constraints, allowing semantic Gaussians to align accurately with planar surfaces while adapting in complex regions. For robust and efficient Gaussian refinement, we introduce a Semantic-aware Uncertainty Sampler that actively selects Gaussians to update memory. Our approach is tailored for embodied scene understanding, preserving planar structures and sharp boundaries during progressive indoor exploration.
@@ 2.3. Indoor 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) and its variants (Szymanowicz et al., 2024; Qian et al., 2024; Lin et al., 2024; Wei et al., 2024b, c; Wang et al., 2024b) have shown promising rendering results and faster performance compared to NeRF (Mildenhall et al., 2021). However, vanilla optimization for 3DGS leads to disorganized Gaussian distributions due to the absence of explicit geometric constraints, resulting in poor surface continuity and imprecise object boundaries. Several approaches address these limitations, including DN-Splatter (Turkulainen et al., 2024) and GSDF (Yu et al., 2024a), which incorporate geometric priors, while SuGaR (Guédon and Lepetit, 2024), PGSR (Chen et al., 2024), and RaDe-GS (Zhang et al., 2024) employ flattened Gaussians to improve surface reconstruction.
@@ 4.1. Settings
Datasets. Our model is evaluated on the EmbodiedOcc-ScanNet dataset (Wu et al., 2024), which is based on the Occ-ScanNet dataset (Yu et al., 2024b). This dataset provides monocular images paired with voxel-level semantic occupancy annotations. The dataset is divided into 537 training scenes and 137 validation scenes, with each scene comprising 30 posed frames and their corresponding occupancy annotations. In addition to the full dataset, EmbodiedOcc-ScanNet also provides mini versions: EmbodiedOcc-ScanNet-mini with 64/16 scenes in train/val splits respectively. The occupancy representation is structured as a $60\times 60\times 36$ voxel grid (corresponding to a $4.8m\times 4.8m\times 2.88m$ space in front of the camera) with a voxel resolution of $0.08m$. For global occupancy representations, the resolution varies based on scene dimensions and is calculated as $(lx\times ly\times lz)/0.08m$, where $lx,ly,lz$ represents the scene’s spatial extent in world coordinates. The semantic annotations in this dataset encompass 12 distinct classes, comprising $11$ meaningful object and structural categories including architectural elements (ceiling, floor, wall, window), furniture (chair, bed, sofa, table), electronic devices (TVs), and general object classifications alongside an additional empty space class.
@@
Figure 4. Visualization of our local occupancy prediction. Our method accurately reconstructs complex geometric structures and preserves fine details in challenging indoor scenes. The visualization demonstrates improved performance at object boundaries and thin structures.
@@ 4.2. Implementation
Parameter Configuration. For our Geometry-guided Refinement module, we set curvature thresholds $\kappa_{low}=5.0$ and $\kappa_{high}=20.0$, normal constraint weights bounded by $w_{min}=0.0$ and $w_{max}=1.0$, distance thresholds $d_{near}=0.1$ and $d_{far}=0.25$, and employ $10$ neighbor points during nearest neighbor search. For our Semantic-aware Uncertainty Sampler used in embodied occupancy prediction, we initialize Gaussians with a $0.16m$ interval for scene representation. During each update, we set Monte Carlo dropout sampling times to $3$ and the uncertainty threshold to $0.3$ in the online refinement layer. We use the same loss functions and weighting scheme as EmbodiedOcc (Wu et al., 2024).
@@ 4.3. Main Results
As shown in Tab. 2, EmbodiedOcc++ consistently outperforms both the original EmbodiedOcc (Wu et al., 2024) and the SplicingOcc baseline across all evaluation settings. Remarkably, these improvements are achieved without introducing any additional trainable parameters, demonstrating the effectiveness of our geometry-aware and uncertainty-driven techniques within the same architectural backbone. Our method exhibits stronger performance across both structural (e.g., wall, furniture) and object-centric (e.g., bed, sofa, table) categories, indicating improved spatial consistency and semantic fidelity. These gains stem from the Geometry-guided Refinement Module, which encourages planar alignment, and the Semantic-aware Uncertainty Sampler, which improves update reliability in overlapping regions. Qualitative results in Fig. 5 demonstrate that our approach yields more consistent and structurally precise occupancy predictions, effectively preserving sharp edges and planar boundaries across sequential updates from multiple viewpoints. For more visualization results, please see our appendix and video in the supplementary materials.
@@ 4.4. Component Analysis
Analysis of the Uncertainty Point Update. We investigate the impact of uncertainty thresholds in our proposed Semantic-aware Uncertainty Sampler on embodied occupancy prediction. Tab. 5 shows that with a high threshold of 0.7, the model exhibits conservative update behavior. Reducing the threshold to 0.5 improves both IoU and mIoU scores. Optimal performance is achieved with a threshold of 0.3, resulting in significantly improved metrics. Additionally, setting the threshold too low (e.g., 0.1) diminishes the effect of uncertainty sampling, leading to redundant updates and degraded performance. This validates our approach of using entropy as an update ratio to determine which regions benefit from new observations. Our uncertainty update improves consistency when handling overlapping regions observed from different viewpoints. These results highlight the importance of this mechanism for robust embodied occupancy prediction in sequential observation scenarios.
@@
4.5. Supplementary Material
For a comprehensive understanding of our framework, we include extended qualitative visualizations, ablation studies, and implementation details in the Appendix. These cover local and embodied occupancy predictions, detailed analysis of normal constraint fusion strategies, and computational efficiency. In addition, we provide a video showcasing progressive occupancy prediction results under various embodied exploration scenarios. Due to the space limitation, please refer to the supplementary materials for more details.

### v2

@@ Abstract.
3D occupancy prediction; 3D Gaussian Splatting; Online scene understanding
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754740isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Scene understanding
Figure 1. EmbodiedOcc++ Framework: A monocular RGB-based indoor 3D occupancy prediction method enhanced with plane regularization and uncertainty sampling. The Geometry-guided Refinement Module regularizes Gaussian updates along surface tangent planes using adaptive curvature and depth constraints, while the Semantic-aware Uncertainty Sampler enables efficient Gaussian selection. Designed for embodied scene understanding with preserved planar structures during progressive indoor exploration.
@@ 2.3. Indoor 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) and its variants (Szymanowicz et al., 2024; Qian et al., 2024; Lin et al., 2024; Wei et al., 2024b, c; Wang et al., 2024b) have shown promising rendering results and faster performance compared to NeRF (Mildenhall et al., 2021). However, vanilla optimization for 3DGS leads to disorganized Gaussian distributions due to the absence of explicit geometric constraints, resulting in poor surface continuity and imprecise object boundaries. Several approaches address these limitations, including DN-Splatter (Turkulainen et al., 2024), GSDF (Yu et al., 2024a) and OmniIndoor3D (Wei et al., 2025), which incorporate geometric priors, while SuGaR (Guédon and Lepetit, 2024), PGSR (Chen et al., 2024), and RaDe-GS (Zhang et al., 2024) employ flattened Gaussians to improve surface reconstruction.
@@ 4.1. Settings
[…]
@@
Figure 4. Visualization of our local occupancy prediction. Our method accurately reconstructs complex geometric structures and preserves fine details in challenging indoor scenes.
@@ 4.2. Implementation
Parameter Configuration. For our Geometry-guided Refinement module, we set curvature thresholds $\kappa_{low}=5.0$ and $\kappa_{high}=20.0$, normal constraint weights bounded by $w_{min}=0.0$ and $w_{max}=1.0$, distance thresholds $d_{near}=0.1$ and $d_{far}=0.25$, and employ 101010 neighbor points during nearest neighbor search. For our Semantic-aware Uncertainty Sampler used in embodied occupancy prediction, we initialize Gaussians with a $0.16m$ interval for scene representation. During each update, we set Monte Carlo dropout sampling times to 333 and the uncertainty threshold to 0.30.30.3 in the online refinement layer. We use the same loss functions and weighting scheme as EmbodiedOcc (Wu et al., 2024).
@@ 4.3. Main Results
As shown in Tab. 2, EmbodiedOcc++ consistently outperforms both the original EmbodiedOcc (Wu et al., 2024) and the SplicingOcc baseline across all evaluation settings. Remarkably, these improvements are achieved without introducing any additional trainable parameters, demonstrating the effectiveness of our geometry-aware and uncertainty-driven techniques within the same architectural backbone. Our method exhibits stronger performance across both structural (e.g., wall, furniture) and object-centric (e.g., bed, sofa, table) categories, indicating improved spatial consistency and semantic fidelity. These gains stem from the Geometry-guided Refinement Module, which encourages planar alignment, and the Semantic-aware Uncertainty Sampler, which improves update reliability in overlapping regions. Qualitative results in Fig. 5 demonstrate that our approach yields more consistent and structurally precise occupancy predictions, effectively preserving sharp edges and planar boundaries across sequential updates from multiple viewpoints.
@@ 4.4. Component Analysis
Analysis of the Uncertainty Point Update. We investigate the impact of uncertainty thresholds in our proposed Semantic-aware Uncertainty Sampler on embodied occupancy prediction. Tab. 5 shows that with a high threshold of 0.7, the model exhibits conservative update behavior. Reducing the threshold to 0.5 improves both IoU and mIoU scores. Optimal performance is achieved with a threshold of 0.3, resulting in significantly improved metrics. Additionally, setting the threshold too low diminishes the effect of uncertainty sampling, leading to redundant updates and degraded performance. This validates our approach of using entropy as an update ratio to determine which regions benefit from new observations. Our uncertainty update improves consistency when handling overlapping regions observed from different viewpoints. These results highlight the importance of this mechanism for robust embodied occupancy prediction in sequential observation scenarios.
@@ 5. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (62476011).